# A Statistical Machine Learning Approach for Adapting Reduced-Order Models using Projected Gaussian Process

### v1

@@ Abstract
Projection-based model reduction is among the most widely adopted approaches for constructing parametric Reduced-Order Models (ROMs). Utilizing snapshot data from solving full-order governing equations, the Proper Orthogonal Decomposition (POD) computes the optimal basis modes that span a low-dimensional subspace where the ROM resides. Because a governing equation is often parameterized by a set of parameters, challenges immediately arise when one would like to investigate how systems behave differently over the parameter space (in design, diagnosis, control, uncertainty quantification and real-time operations). In this case, the optimal basis needs to be efficiently updated so as to adapt ROM that can accurately capture the variation of a system’s behavior over its parameter space. In this paper, we propose a Projected Gaussian Process (pGP) model and formulate the problem of adapting POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping (injective) from the parameter space to the Grassmann Manifold that contains the optimal vector subspaces. To establish such a relationship, a mapping is found between the Euclidean space and the horizontal space of an orthogonal matrix that spans a reference subspace in the Grassmann Manifold. Then, a second mapping from the horizontal space to the Grassmann Manifold is established through the Exponential/Logarithm maps between the manifold and its tangent space. Finally, given a new parameter, the conditional distribution of a vector can be found in the Euclidean space using the Gaussian Process (GP) regression, and such a distribution is then projected to the Grassmann Manifold that enables us to find the optimal subspace, i.e., POD basis, for the new parameter. Compared with existing interpolation method, the proposed statistical learning approach allows us to optimally estimate (or tune) model parameters given data (i.e., the prediction/interpolation becomes problem-specific), and quantify the uncertainty associated with the prediction. Numerical examples are presented to demonstrate the advantages of the proposed pGP for adapting POD basis against parameter changes.
keywords:
Reduced-Order Models, Proper Orthogonal Decomposition, Gaussian Process, Projected Gaussian Process, Grassmann Manifold
@@ 1.1 Background and Problem Statement
Consider the diffusion of chemical species, aeroelasticity of aircraft wings, flow of vehicles and predator-prey interactions, dynamics of such complex systems is typically described by a set of Partial Differential Equations (PDEs). Very often, the dimensions of these problems are extremely high due to the spatial discretization of the mathematical models (e.g., Computational Fluid Dynamics models or Finite Element Analysis with very large numbers of degrees of freedom). The prohibitive computational cost prevents us from repeatedly running numerical solvers under different parameter settings; often needed for design, diagnosis, prediction, uncertainty quantification and control problems[24, 25, 29]. Hence, when there exist low-dimensional patterns embedded in high-dimensional systems[16], Reduced-Order Models (ROMs) become indispensable in capturing the dominate behaviors of complex systems. Compared with data-driven surrogate models or meta models[3], ROM often preserve important system physics, such as stability[26, 4], passivity[31, 9], polynomial nonlinearities[27], etc.
Projection-based model reduction is among the most widely adopted approaches for constructing ROMs[16, 27]. The central idea behind projection-based ROM is to project a high-dimensional governing equation to an optimal low-dimensional vector space spanned by orthogonal bases. Although searching for such bases may seem to be nontrivial considering the computational speed, accuracy and complex boundary conditions, the Proper Orthogonal Decomposition (POD)—applied to curves in Hilbert spaces $\mathcal{H}$ of infinite dimension—provides a computationally efficient method for approximating data generated from full-order systems in an $\ell_{2}$ sense[6]. This can be efficiently done by performing the thin Singular Value Decomposition (SVD) of the snapshot data matrix[16].
However, a governing equation is often parameterized by a set of parameters, and challenges immediately arise when one would like to investigate how systems behave differently over the parameter space (of the governing equation). In this case, the optimal basis needs to be updated so that the ROM can accurately capture the variation of a system’s behavior over its parameter space. As pointed out by [10, 32], the robustness and accuracy of an ROM are related to the reference parameters from which POD modes are derived. The accuracy of ROM determines the performance of any downstream machine learning models built on the low-dimensional ROM.
@@
(1) $(t,\bm{s})\in[0,T]\times\mathbb{S}\mapsto x(t,\bm{s};\bm{\theta})$
where $\mathbb{S}$ is the spatial domain and $T>0$. Let $\bm{x}(t;\bm{\theta})=(x(t,\bm{s}_{1};\bm{\theta}),x(t,\bm{s}_{2};\bm{\theta}),\cdots,x(t,\bm{s}_{n};\bm{\theta}))$ be an $n$-dimensional vector representing the discretized solutions at a set of locations $\{\bm{s}_{1},\bm{s}_{2},\cdots,\bm{s}_{n}\}\in\mathbb{R}^{n}$ (say, using the Finite Element method), one can obtain the so-called snapshot data matrix, $\bm{D}(\bm{\theta})=[\bm{x}(t_{1};\bm{\theta});\bm{x}(t_{2};\bm{\theta});\cdots;\bm{x}(t_{n_{T}};\bm{\theta})]$, where $n_{T}$ is the number of temporal points. The POD method finds the optimal $r$-dimensional ($r\ll n$) vector subspace $\mathbb{W}\subset\mathbb{R}^{n}$ and an orthogonal projection, $\mathcal{W}:\mathbb{R}^{n}\rightarrow\mathbb{W}$ such that the following distance is minimized[12]
(2) $\text{min}_{\mathcal{W}}\sum_{i=1}^{n_{T}}||\bm{x}(t_{i};\bm{\theta})-\mathcal{W}(\bm{x}(t_{i};\bm{\theta}))||^{2}$
@@
Hence, we see both the advantages and limitations of the POD approach above. The optimal POD basis, which is obtained from the snapshot data generated at the parameter setting $\bm{\theta}$, is optimal and only optimal for parameter $\bm{\theta}$. As shown in Figure 1, if data are already available from numerically solving the full-order governing equation repeatedly at a set of pre-determined parameter settings, it is possible to obtain the optimal basis for each parameter setting using the existing approach. However, for any new parameter in the parameter space, new basis modes (which is optimal or near optimal for the new parameter) is needed so that one can adapt the ROM according to parameter changes. The lack of robustness against parameter change limits the applicability of ROMs for design, diagnosis, control, uncertainty quantification and real-time operations (e.g., digital twins), where parameter changes are common[1, 24, 25, 29].
In this paper, we formulate the problem of adapting POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping between the parameter $\bm{\theta}$ and POD basis $\bm{\Phi}$ that spans a low-dimensional vector subspace. This subspace regression problem can be stated as follows:
@@
We address this problem by proposing a supervised statistical learning approach, known as the Projected Gaussian Process (pGP) regression. As shown in Figure 1, when (training) data are available by numerically solving the full-order governing equation repeatedly at a (usually small) number of parameter settings $\{\bm{\theta}_{i}\}_{i=1}^{k}$, it is possible to obtain the optimal basis $\{\bm{\Phi}_{i}\}_{i=1}^{k}$ at each parameter setting, separately, using the existing approach. Once the proposed pGP regression learns the mapping between parameters and POD bases, a new basis can be generated given the new parameters, thus adapting the ROM for parameter changes (see Sec.3).
Figure 1: A supervised statistical learning problem (based on the proposed Projected Gaussian Process) that establishes the mapping between parameters and POD basis modes so that ROMs can be adapted for parameter changes.
@@ 1.2 Literature Review
Various strategies have been proposed to adapt ROM for parameter changes. Figure 2 presents one possible way to group existing approaches into four categories. This figure is used to facilitate the explanations of different approaches as well as how the proposed approach is built upon prior work. The boundaries that separate different approaches may change if we examine these methods from different perspectives.
@@
$\bullet$ Strategy I involves pre-computing a database, i.e., a library, of reduced-order bases[1, 15]. For applications where this approach is feasible, ROM can be updated in real time over the set of parameters included in the database.
@@
$\bullet$ Strategy II computes the first-order total derivatives of the POD modes with respective to a parameter[8], and a new POD basis, given a small change of the parameters from a base point, can be found through the first-order expansion in the parametric space.
@@
$\bullet$ Strategy III involves different approaches that aim to interpolate or predict new POD bases from the observed parameter-POD basis pairs. In particular, the POD basis interpolation on Grassmann Manifolds have been extensively investigated[2, 30, 6, 23]. This approach is based on an important result that a subspace spanned by a POD basis is a point of the Grassmann Manifold that consists of $r$-dimensional subspaces in $\mathbb{R}^{n}$, where $n$ and $r$ are respectively the original and reduced-order dimensions. Hence, a tangent space, attached to a given base point in the Grassmann Manifold, can be constructed, and the subspaces from the Grassmann Manifold can be projected to the tangent space through a logarithmic map. After this critical step, interpolation can be done on the tangent space (such as the Lagrangian interpolation[2], inverse distance weighting[23], or Gaussian Process[13]), and the interpolated vector on the tangent space is eventually mapped back to the Grassmann Manifold through an exponential map. Hence, this method depends on the choice of the base point upon which the tangent space is attached to, and naturally requires the injective radius of the exponential map to be sufficiently large so that the approach is stable[12]. In contrast, the subspace angle interpolation approach directly computes the geometric distance between two subspaces by finding the principal angels between two POD bases[34]. The principal angels can be viewed as a series of angles required to rotate one subspace to form another, and such a distance metric is used to interpolate subspaces.
@@
$\bullet$ Strategy IV involves statistical machine learning to learn the mapping between parameters and POD bases (or the subspaces spanned by POD bases). Leveraging the key ideas from both the POD basis interpolation on Grassmann Manifold[2] and subspace angle interpolation[34], a regression tree-based approach on Grassmann Manifold was proposed in [18]. Following this approach, a tree is grown by repeatedly splitting the tree node (i.e., a binary partition the parameter space) to maximize the Riemannian distance between two subspaces (based on the principals angles) spanned by the predicted POD bases on the left and right daughter nodes. As a result, this tree-based approach effectively divides the parameter space into sub-regions, and find the optimal POD basis for each sub-region (which could generate useful insights on how a system’s behavior varies over the parameter space). This approach successfully combines the POD basis prediction problem into the classical framework of regression trees, and appears to be more stable and conservative than the existing POD basis interpolation on Grassmann Manifolds (as the tree-based method only updates the POD basis when the parameter jumps from one sub-region to another in the parameter space). We also note that, this approach is useful for finding the POD basis when the dimension of the parameter space is high and the system behaviors have different sensitivity levels against different parameters. For those less important parameters, the tree-based approach effectively exclude those parameters from being used to split the tree (note that, the existing interpolation method does not automatically do this).
@@
Other statistical learning approaches involve the use of Gaussian Process (GP). [13] directly used GP to capture the relationship between parameters and the matrices that determine the vectors on the tangent space of the Grassmann Manifold. Hence, given a new parameter setting, GP is used to predict a new vector in the tangent space, which yields the predicted subspace on the Grassmann Manifold through an exponential map. Note that, because GP performs prediction through a weighted linear combination of responses, this approach is similar to the existing Lagrangian POD basis interpolation on Grassmann Manifolds but essentially with different interpolation weights. For the reasons noted in Sec.3, we refer this approach as the native GP approach following the terminology in [21]. [20] proposed an approach known as Common POD (cPOD). In the context of engine injector design, a common POD basis is found for some base design geometry, the POD bases for other geometries are obtained by rescaling the common POD through a linear map. This approach requires sound domain knowledge of the system under consideration to determine the base design and establish the linear map between POD bases from the common POD. [35] proposed a GP subspace regression where the central idea is to directly construct a matrix-variate GP for the matrix bases such that the approach does not hinge on the differential geometric structures of the Gassmann manifold.
@@ 1.3 Contributions of this Paper
The prior work in the literature has shown the great potential of leveraging machine learning for adapting projection-based ROMs. The main contribution of this paper is to propose a new supervised statistical learning method, known as the Projected Gaussian Process (pGP) regression, that enables the learning of the mapping between parameters and POD bases. To our best knowledge, this method is not yet available in the literature. Computer code is also available as Supplementary Materials.
@@
As a statistical machine learning approach, the proposed pGP provides some critical capabilities much needed for adapting ROMs against parameter changes:
@@
i). Unlike existing interpolation method (e.g., Lagrangian interpolation), the statistical learning approach allows us to optimally estimate (or tune) model parameters based on the training data (i.e., the prediction becomes problem-specific). In contrast, when the Lagrangian interpolation is used, the Lagrange bases are computed from pre-existing equations which are independent of the problem or data of interest. Because the pGP model is always trained and optimized given the training dataset, such a flexibility potentially provides the opportunity to further improve the accuracy of ROM for the specific problem at hand (as shown in the numerical examples)
ii). The proposed pGP, as a statistical learning approach, provides the capability of parameter selection when constructing the mapping between parameters and POD bases. This capability is largely missing in the literature but is particularly important for systems with high-dimensional parameter space, i.e., not all parameters are equally important in terms of how they change the POD basis modes. If relevant parameters can be identified, one only needs to explore a reduced parameter space, with a lower dimension, when investigating how system dynamics varies in the parameter space.
iii). Due to the parameter estimation/selection capabilities, the pGP approach naturally involves the use of global POD basis as its special case. When no parameter is identified as significant, the predicted POD basis matrices for different parameters, under the proposed pGP framework, become i.i.d. samples from a matrix-variate Gaussian distribution with the global POD basis matrix being the mean of this distribution (as shown in numerical Example II).
@@
iv). The proposed pGP not only provides the point prediction of the POD basis, but also scientifically quantifies the uncertainty associated with the predicted POD basis as well as the output generated by ROMs. Such a capability of uncertainty quantification facilitates decision-making under uncertainty, and provides a critical link (as a future research path) from the proposed approach to a large body of literature on statistical experimental design and deep reinforcement learning for more efficient construction of ROMs by sequentially exploring the solution space of full-order models[14, 7].
@@
The paper is organized as follows. In Section 2, we provide the preliminaries on POD basis interpolation. The proposed pGP approach is presented in Section 3. Comprehensive numerical examples and discussions are presented in Section 4. Section 5 concludes the paper.

### v2

# Adapting Projection-Based Reduced-Order Models using Projected Gaussian Process

@@ Abstract
Projection-based model reduction is among the most widely adopted methods for constructing parametric Reduced-Order Models (ROM). Utilizing the snapshot data from solving full-order governing equations, the Proper Orthogonal Decomposition (POD) computes the optimal basis modes that represent the data, and a ROM can be constructed in the low-dimensional vector subspace spanned by the POD basis. For parametric governing equations, a potential challenge arises when there is a need to update the POD basis to adapt ROM that accurately capture the variation of a system’s behavior over its parameter space (in design, control, uncertainty quantification, digital twins applications, etc.). In this paper, we propose a Projected Gaussian Process (pGP) and formulate the problem of adapting the POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping from the parameter space to the Grassmann manifold that contains the optimal subspaces. A mapping is firstly established between the Euclidean space and the horizontal space of an orthogonal matrix that spans a reference subspace in the Grassmann manifold. A second mapping from the horizontal space to the Grassmann manifold is established through the Exponential/Logarithm maps between the manifold and its tangent space. Finally, given a new parameter, the conditional distribution of a vector can be found in the Euclidean space using the Gaussian Process (GP) regression, and such a distribution is then projected to the Grassmann manifold that enables us to predict the optimal subspace for the new parameter. As a statistical learning approach, the proposed pGP allows us to optimally estimate (or tune) the model parameters from data and quantify the statistical uncertainty associated with the prediction. The advantages of the proposed pGP are demonstrated by numerical experiments.
Keywords: Reduced-Order Models, Grassmann manifold, Proper Orthogonal Decomposition, Gaussian Process, Projected Gaussian Process.
@@ 1.1 Background and Problem Statement
Consider the diffusion of chemical species, aeroelasticity of aircraft wings, flow of vehicles and predator-prey interactions, dynamics of such complex systems is typically described by a set of Partial Differential Equations (PDEs). Very often, the dimensions of these problems are high after the spatial discretization of the mathematical models (e.g., computational fluid dynamics models or finite element analysis with very large numbers of degrees of freedom). The prohibitive computational cost prevents us from repeatedly running numerical solvers under different parameter settings; often needed for design, diagnosis, prediction, uncertainty quantification and digital twins applications (Peherstorfer and Willcox, 2015a, ; Peherstorfer and Willcox, 2015b, ; Sargsyan et al.,, 2015). Hence, when there exist low-dimensional patterns embedded in high-dimensional systems (Kutz et al.,, 2016), Reduced-Order Models (ROM) become indispensable in capturing the dominate behaviors of complex systems. Compared with data-driven surrogate models or meta models (Barton and Meckesheimer,, 2006), ROM often preserve important system physics, such as stability (Prajna,, 2003; Batool et al.,, 2021), passivity (Sorensen,, 2003; Breiten and Unger,, 2022), polynomial nonlinearities (Qian et al.,, 2020), etc.
Among various model reduction approaches, projection-based model reduction is perhaps the most widely adopted one for constructing ROM (Kutz et al.,, 2016; Qian et al.,, 2020). The central idea behind projection-based ROM is to project a high-dimensional system to an optimal low-dimensional vector space spanned by orthogonal bases. Although searching for such bases may seem to be a nontrivial task considering the computational speed, accuracy and complex boundary conditions, the Proper Orthogonal Decomposition (POD)—applied to curves in Hilbert spaces $\mathcal{H}$ of infinite dimension—provides a computationally efficient approach that optimally approximates data generated from full-order systems in an $\ell_{2}$ sense (Berkooz et al.,, 1993; Rowley,, 2005; Benner et al.,, 2015). This can be efficiently done by performing the thin Singular Value Decomposition (SVD) of the snapshot data matrix, and one may refer to Kutz et al., (2016) for a more detailed description of the procedure.
@@
$(t,\bm{s})\in[0,T]\times\mathbb{S}\mapsto x(t,\bm{s};\bm{\theta})$ (1)
where $\mathbb{S}$ is the spatial domain and $T>0$. Let $\bm{x}(t;\bm{\theta})=(x(t,\bm{s}_{1};\bm{\theta}),x(t,\bm{s}_{2};\bm{\theta}),\cdots,x(t,\bm{s}_{n};\bm{\theta}))$ be an $n$-dimensional vector representing the discretized solutions at a set of locations $\{\bm{s}_{1},\bm{s}_{2},\cdots,\bm{s}_{n}\}$, one obtains the snapshot data matrix, $\bm{D}(\bm{\theta})=[\bm{x}(t_{1};\bm{\theta});\bm{x}(t_{2};\bm{\theta});\cdots;\bm{x}(t_{n_{T}};\bm{\theta})]$ with $n_{T}$ being the number of temporal points. The POD method finds the optimal $r$-dimensional ($r\ll n$) vector subspace $\mathbb{W}\subset\mathbb{R}^{n}$ and an orthogonal projection, $\mathcal{W}:\mathbb{R}^{n}\rightarrow\mathbb{W}$, such that the following distance is minimized
@@
where $||\cdot||=\sqrt{\left\langle\cdot,\cdot\right\rangle}$ and $\left\langle\cdot,\cdot\right\rangle$ is the inner product of the Hilbert space.
[…]
Hence, for parametric dynamical systems, it is critical for the POD modes to be sufficiently representative so that data generated from various parameter settings are well represented. This is particularly important for design, diagnosis, control, uncertainty quantification and real-time operations (e.g., digital twins) where parameter changes are common (Amsallem et al.,, 2009; Peherstorfer and Willcox, 2015a, ; Peherstorfer and Willcox, 2015b, ; Sargsyan et al.,, 2015). One approach, known as the greedy basis generation, finds the reduced basis by an iterative procedure (Hesthaven et al.,, 2016). At each iteration, one new basis function is added using one solution computed at a new parameter setting. Throughout multiple iterations, the overall precision is gradually enhanced and the final basis is rich enough to represent data from the parameter space.
Alternatively, the optimal basis can be adapted as parameters change so that the POD basis can capture the variation of a system’s behavior over its parameter space (Bui-Thanh et al.,, 2008; Vetrano et al.,, 2011). In this paper, we focus on this approach and formulate the problem of adapting the POD basis as a supervised statistical learning problem. The goal is to learn a mapping between the parameter $\bm{\theta}$ and POD basis $\bm{\Phi}$ that spans a low-dimensional vector subspace. Such a subspace regression problem is stated as follows:
@@
We address this problem by proposing a supervised statistical learning approach, known as the Projected Gaussian Process (pGP) regression. As shown in Figure 1, when (training) data are available by numerically solving the full-order governing equation repeatedly at a (usually small) number of parameter settings $\{\bm{\theta}_{i}\}_{i=1}^{k}$, it is possible to obtain the optimal basis $\{\bm{\Phi}_{i}\}_{i=1}^{k}$ at each parameter setting, separately, using the existing approach. Once the proposed pGP regression learns the mapping between parameters and POD bases, a new basis can be predicted given the new parameters, thus adapting the ROM for parameters change (see Sec.3).
Figure 1: A supervised statistical learning problem based on the proposed pGP that enables the prediction of POD basis modes as parameters change.
@@ 1.2 Literature Review
Various approaches have been investigated for tackling the model reduction problem of parametric dynamical systems. Some examples include the interpolatory model reduction and rational interpolation (Baur et al.,, 2011; Gosea et al.,, 2021), multi-moment matching (Freund,, 2003; Benner and Feng,, 2014), balanced truncation (Wittmuess et al.,, 2016), matrix interpolation (Panzer et al.,, 2010; Geuss et al.,, 2013), etc. For a comprehensive and complete review of projection-based model reduction for parametric dynamical systems, we refer the readers to (Benner et al.,, 2015, 2017).
@@
In this section, we primarily focus on major strategies for adapting the POD basis for parameters change. Figure 2 presents one possible way to group existing approaches into four categories. This figure is only used to facilitate the explanations of different approaches as well as how the proposed approach is built upon the prior work. The boundaries that separate different approaches may change if we examine these methods from different perspectives.
@@
$\bullet$ Strategy I involves pre-computing a database, i.e., a library, of reduced-order bases (Amsallem et al.,, 2009; Kapteyn et al.,, 2020). For applications where this approach is feasible, ROM can be updated in real time over the set of parameters included in the database.
@@
$\bullet$ Strategy II computes the first-order total derivatives of the POD modes with respective to a parameter (Borggaard and Pelletier,, 2008), and a new POD basis, given a small change of the parameters from a base point, can be found through the first-order expansion in the parametric space.
@@
$\bullet$ Strategy III involves different approaches that aim to interpolate or predict new POD bases from the observed parameter-POD basis pairs. In particular, the POD basis interpolation on Grassmann manifolds have been extensively investigated (Amsallem and Farhat,, 2008; Son,, 2013; Benner et al.,, 2015; Mosquera et al.,, 2019). This approach is based on an important result that a subspace spanned by a POD basis is a point on the Grassmann manifold that consists of $r$-dimensional subspaces in $\mathbb{R}^{n}$, where $n$ and $r$ are respectively the original and reduced-order dimensions. Hence, a tangent space, attached to a given base point on the Grassmann manifold, can be constructed, and the subspaces from the Grassmann manifold can be projected to the tangent space through a logarithmic map. After this critical step, interpolation can be done on the tangent space (such as the Lagrangian interpolation (Amsallem and Farhat,, 2008), inverse distance weighting (Mosquera et al.,, 2019), or Gaussian Process (Giovanis and Shields,, 2020)), and the interpolated vector on the tangent space is eventually mapped back to the Grassmann manifold through an exponential map. Hence, this method depends on the choice of the base point upon which the tangent space is attached to, and naturally requires the injective radius of the exponential map to be sufficiently large so that the approach is stable (Friderikos et al.,, 2020). In contrast, the subspace angle interpolation approach directly computes the geometric distance between two subspaces by finding the principal angels between two POD bases (Ye and Lim,, 2016). The principal angels can be viewed as a series of angles required to rotate one subspace to form another, and such a distance metric is used to interpolate subspaces. Other examples based on the idea of interpolation include the matrices interpolation of local reduced models for linear time-invariant systems (Panzer et al.,, 2010), and Son, (2013) investigated the interpolation on Grassmann manifolds for affinely dependent system matrices on model parameters.
@@
$\bullet$ Strategy IV involves using statistical machine learning to learn the mapping between parameters and POD bases (or the subspaces spanned by POD bases). Leveraging the key ideas from both the POD basis interpolation on Grassmann manifold (Amsallem and Farhat,, 2008) and subspace angle interpolation (Ye and Lim,, 2016), a regression tree-based approach on Grassmann manifold was proposed in Liu and Liu, (2023). Following this approach, a tree is grown by repeatedly splitting the tree node (i.e., a binary partition the parameter space) to maximize the Riemannian distance between two subspaces (based on the principals angles) spanned by the predicted POD bases on the left and right daughter nodes. As a result, this tree-based approach effectively divides the parameter space into sub-regions, and the POD basis for a sub-region is obtained using only the data generated from that sub-region. This approach successfully combines the POD basis prediction problem into the classical framework of regression trees, and appears to be more stable and conservative than the existing POD basis interpolation on Grassmann manifolds (as the tree-based method only updates the POD basis when the parameter jumps from one sub-region to another in the parameter space). We also note that, this approach is useful for finding the POD basis when the dimension of the parameter space is high and the system behaviors have different sensitivity levels against different parameters. For those less important parameters, the tree-based approach effectively exclude those parameters from being used to split the tree (note that, the existing interpolation method does not automatically do this). Finally, in very recent years, neural networks have also been used for adapting ROM. For example, Franco et al., (2024) proposed the Deep Orthogonal Decomposition by constructing a deep neural network to approximate the solution manifold through a continuously adaptive local basis (i.e., use a neural network to learn the mapping from parameters and subspaces). Berman and Peherstorfer, (2024) proposed to continuously adapt the low-rank weights of pre-train neural networks to predict the evolution of solution fields at new physics parameters. In general, training a neural network requires a large amount of training data (e.g., Franco et al., (2024) used 500 training samples for model reduction of the Eikonal equation). This may potentially pose additional challenges when it is expensive to generate training data by solving full-order physics.
@@
Other statistical learning approaches involve the use of GP. For example, Giovanis and Shields, (2020) directly used GP to capture the relationship between parameters and the matrices that determine the vectors on the tangent space of the Grassmann manifold. Hence, given a new parameter setting, GP is used to predict a new vector in the tangent space, which yields the predicted subspace on the Grassmann manifold through an exponential map. Note that, because GP performs prediction through a weighted linear combination of responses, this approach is similar to the existing Lagrangian POD basis interpolation on Grassmann manifolds but essentially with different interpolation weights. For the reasons noted in Sec.3, we refer to this approach as the native GP approach following the terminology in Mallasto and Feragen, (2018). Mak et al., (2018) proposed another approach known as the Common POD (cPOD). In an engine injector design example, a common POD basis is found for some base design geometry, the POD bases for other geometries are obtained by rescaling the common POD through a linear map. This approach requires sound domain knowledge of the system under consideration to determine the base design and establish the linear map between POD bases from the common POD. Zhang et al., (2022) proposed a GP subspace regression where the central idea is to directly construct a matrix-variate GP for the matrix bases such that the approach does not hinge on the differential geometric structures of the Gassmann manifold.
@@ 1.3 Contributions of this Paper
The prior work in the literature has shown the great potential of leveraging machine learning for adapting projection-based ROM. The main contribution of this paper is to propose a new supervised statistical learning method that enables the learning of the mapping between parameters and POD bases. To our best knowledge, this pGP approach is not yet available in the literature.
@@
As a statistical machine learning approach, the proposed pGP provides some critical capabilities much needed for adapting ROM against parameter changes:
@@
i). Unlike existing interpolation methods (e.g., Lagrangian interpolation), the proposed statistical learning approach allows us to optimally estimate (or tune) model parameters based on training data (i.e., the predictions are data-driven). In contrast, when the Lagrangian interpolation is used, the Lagrange bases are computed from pre-existing equations which are independent of the problem or data of interest. Because the pGP model is always trained and optimized given a training dataset, such a flexibility provides the opportunity to further improve the accuracy of the predicted POD basis modes for specific problems at hand (as shown in the numerical examples)
@@
ii). The proposed pGP not only provides the point prediction of the POD basis, but also quantifies the statistical prediction uncertainty associated with the predicted POD modes. Such a capability of uncertainty quantification facilitates decision-making under uncertainty, and provides a critical link (as a future research path) to a large body of literature on statistical experimental design and deep reinforcement learning for more efficient construction of the POD basis by sequentially exploring the solution space of full-order models (Ivanova et al.,, 2021; Blau et al.,, 2022).
iii). The proposed pGP, as a statistical learning approach, provides the capability of parameter selection when constructing the mapping between parameters and POD bases. This capability is needed when not all parameters are equally important in terms of their predictive power for POD modes. If important parameters can be identified, the statistical pGP model may require a smaller number of input. When experimental design is used to determine the optimal parameter settings for subsequent experiment, this can also potentially lower the dimension of the parameter space (i.e., the decision space in experimental design becomes smaller). We would also like to point out that this argument shares some similarities with the concept of active subspace (Constantine et al.,, 2014). The active subspace is usually a low-dimensional space of physics model input which contains the detected directions of the strongest variability in model output, while the parameter selection mentioned in this paper refers to the process of identifying the parameters which have stronger predictive power of the POD modes in a statistical model. Hence, the outcome of identifying the active subspace allows one to construct an approximation of the physics model on a low-dimensional subspace, while the outcome of parameter selection in this paper allows one to predict the POD modes using a smaller set of parameters.
iii). Due to the parameter estimation/selection capabilities above, the pGP approach naturally involves the use of global POD basis as its special case. As shown in this paper, it is possible to set the parameters in the proposed pGP regression model such that the predicted POD basis matrices for different parameters become i.i.d. samples from a matrix-variate Gaussian distribution with the global POD basis matrix being the mean of this distribution (as shown in the numerical examples).
@@
The paper is organized as follows. In Section 2, we provide the preliminaries on POD basis interpolation. The proposed pGP approach is presented in Section 3. Numerical examples and discussions are presented in Section 4. Section 5 concludes the paper.